if third term of GP is 1 and seventh term is 625 then common ratio is? 

Grade:11

1 Answers

987456321
6 years ago
the general formula for term t in a GP is
Tn=arn-1
where a is the first term
now T7/T3=r6/r2
which implies
625=r4
r=5